# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 9.5. Totally-nested tetra-separations from good bags
Assume first that $o(B)$ or $o(D)$ has a neighbour in $T$ with a triangle-torso in $\mathcal{O}$. Say this is a neighbour $t$ of $o(B)$. Since $y^{W}_{1}$ is the tip of the triangle $H_{o(B)}$, the vertex $y^{W}_{1}$ is an end of the free edge $f$ of $H_{t}$ by Lemma 9.15. Let $u$ denote the other end of $f$. Let $v$ denote the tip of $H_{t}$. Note that $\{u,v\}$ is an adhesion-set, so $x^{W}_{1}\in\{u,v\}$. Hence $x^{W}_{1}$ lies in the two triangle-torsos from the potter corner $AD$ and in the triangle-torso $H_{t}$. Thus $o(D)$ and $t$ are neighbours in $T$ by Lemma 9.15. In particular, $T$ has exactly three nodes. Since $v$ lies in the two triangle-torsos $H_{t}$ and $H_{o(B)}$ while $x^{W}_{1}$ lies in the two triangle-torsos for the potter corner $AD$, we cannot have $v=x^{W}_{1}$ by Lemma 9.15. Hence $x^{W}_{1}=u$. By Lemma 9.4, the two adhesion-sets of $H_{t}$ are spanned by edges in $G$. Hence $x^{W}_{1}y^{W}_{1}$ and $x^{W}_{1}vy^{W}_{1}$ are desired paths.
Figure 9.7. Case $i\geqslant j$ in the proof of Lemma 9.29. Left: $H_{t}$ is a triangle.
@@
Assume otherwise that neither $o(B)$ nor $o(D)$ has a neighbour in $T$ with a triangle-torso. Then the torsos neighbouring $o(B)$ and $o(D)$ via $T$ are 3-connected. Since 3-connected torsos have disjoint adhesion-sets and $i\geqslant j$, we have $|T|=3$. Let $t$ denote the internal node of $T$. Let $\{x^{W}_{1},u\}$ denote the adhesion-set of the edge $o(D)t$, and let $\{y^{W}_{1},v\}$ denote the adhesion-set of the edge $o(B)t$. We find two independent $y^{W}_{1}$$x^{W}_{1}$ paths in $G_{t}$, and are done, unless $H_{t}$ is a 4-wheel with rim $x^{W}_{1}uy^{W}_{1}v$ and $x^{W}_{1}u,y^{W}_{1}v\notin E(G)$, by Lemma 9.28. But in the latter case, $u$ and $v$ have degree three in $G$, contradicting 4-connectivity. ∎
@@
Say $i=j=1$. If $x_{2}$ or $y_{2}$ is an edge, then we are done by (6.4.1), so assume that both are vertices. By Lemma 9.25, there is an $x^{\prime}_{1}$$y^{\prime}_{1}$ path $P^{\prime}$ in $G_{t(0)}-\{x^{\prime}_{2},y^{\prime}_{2}\}$. Then $P^{\prime}$ contains a $\hat{x}_{1}$$\hat{y}_{1}$ path $P$ that avoids the vertices $x_{2}$ and $y_{2}$, by Lemma 9.26. We find two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$ using Lemma 9.29. Hence $\{x_{1},y_{1}\}$ is 3-linked around $S(U,W)$ by (6.4.3). ∎
@@
If $y_{1}$ or $y_{2}$ is an edge, we are done, so assume that both are vertices. There exist two independent $x^{W}_{1}$$x^{W}_{2}$ paths in $G[W]-\{y_{1},y_{2}\}$ by the (Three-Paths Lemma). (9.21). So it remains to find an $x^{\prime}_{1}$$x^{\prime}_{2}$ path in $G^{\prime}:=G_{t_{0}}-\{y^{\prime}_{1},y^{\prime}_{2}\}$ (as any such path contains a suitable $\hat{x}_{1}$$\hat{x}_{2}$ path by Lemma 9.26). Assume for a contradiction that such a path is missing. Then $H_{t(0)}$ is a $K_{4}$ by Lemma 9.17. Since both $y_{1}$ and $y_{2}$ are vertices, the torso $H_{t(1)}$ is 3-connected by Lemma 9.14. Since $H_{t(0)}$ is good, it follows that $H_{t(-1)}$ is a 4-cycle. So $x_{1}$ and $x_{2}$ are the two free edges on this 4-cycle. By 4-connectivity of $G$, both $x^{\prime}_{1}$ and $x^{\prime}_{2}$ have degree at least four. Since $H_{t(-1)}$ is a 4-cycle and $H_{t(0)}$ is a $K_{4}$, it follows that the edge $x^{\prime}_{1}x^{\prime}_{2}$ must exist in $G$, contradicting that there is no $x^{\prime}_{1}$$x^{\prime}_{2}$ path in $G^{\prime}:=G_{t_{0}}-\{y^{\prime}_{1},y^{\prime}_{2}\}$. ∎
@@
Suppose first that $\pi_{1}=\{x_{1},x_{2}\}$ and $\pi_{2}=\{y_{1},y_{2}\}$. We find two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{1},P_{2}$ in $G[W]$ using Lemmas 9.16 and 9.14. The torso $H_{t(0)}$ is 3-connected by assumption, and hence it contains three independent $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths by (Menger for independent paths). (6.2). None of these paths use torso-edges, so all three paths are subgraphs of the bag $G_{t(0)}$. The three $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths can be shortened to three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $Q_{1},Q_{2},Q_{3}$ by Lemma 9.26. Hence $\pi_{1}$ is 0-potter-linked to $\pi_{2}$.
Suppose otherwise that $\pi_{1}=\{x_{1},y_{1}\}$ and $\pi_{2}=\{x_{2},y_{2}\}$, say. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{1},P_{2},P_{3}$ in $G[W]-x^{\prime}_{1}x^{\prime}_{2}-y^{\prime}_{1}y^{\prime}_{2}$ using the (Three-Paths Lemma). (9.21). By Lemma 9.28, we find two independent $\{x^{\prime}_{1},y^{\prime}_{1}\}$$\{x^{\prime}_{2},y^{\prime}_{2}\}$ paths in the bag $G_{t(0)}$, which shorten to two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{4},P_{5}$ in $G_{t(0)}$ by Lemma 9.26, unless $H_{t(0)}$ is a 4-wheel and both torso-edges $x^{\prime}_{1}x^{\prime}_{2},y^{\prime}_{1}y^{\prime}_{2}$ are missing in $G$. But the latter is impossible: since $H_{t(0)}$ is good, some neighbouring torso would be a triangle or a 4-cycle, but then a vertex in the adhesion-set would have degree three in $G$, contradicting 4-connectivity. ∎
@@
(i). Since $H_{t(0)}$ is 3-connected, the vertex $x^{\prime}_{1}$ has $\geqslant 3$ neighbours in $H_{t(0)}$. Let $u$ be a neighbour of $x^{\prime}_{1}$ in $H_{t(0)}$ distinct from $x^{\prime}_{2}$ and the end of the edge $x_{1}$ besides $x^{\prime}_{1}$. All interior vertices of $H_{t(0)}$ are contained in $U\smallsetminus W$, so $u\in\{y^{\prime}_{1},y^{\prime}_{2}\}$. Say $u=y^{\prime}_{1}$. Assume for a contradiction that $y^{W}_{1}\notin H_{t(0)}$. Then $x^{\prime}_{1}y^{\prime}_{1}$ is a $(W\smallsetminus U)$$(U\smallsetminus W)$ edge, so $(U,W)$ violates the matching-condition, a contradiction.
(ii). Assume for a contradiction that $\{x_{1},x_{2}\}$ is $W$-weird. Then $x_{2}=x^{\prime}_{2}$ since $x_{1}$ is an edge by assumption. Moreover, $x^{\prime}_{2}$ has exactly one neighbour outside $H_{t(0)}$ by weirdness. But then $x_{2}$ should be an edge in $S(U,W)$ by the definition of right-left-shift of $(U^{\prime},W^{\prime})$, contradicting that $x_{2}$ is a vertex. ∎
@@
Assume otherwise, say $y_{1}$ is an inward edge. Then by Lemma 9.33 we have $y^{W}_{1}=y^{\prime}_{1}$ and some $x^{\prime}_{j}$ is a neighbour of $y^{W}_{1}$. Then $j\neq 1$ since otherwise $x^{\prime}_{1}y^{W}_{1}$ is a $(U\smallsetminus W)$$(W\smallsetminus U)$ edge, contradicting the matching-condition. So $j=2$. Then $y_{1}^{\prime}=y_{1}^{W}$ is the only neighbour of $x_{2}$ in $W\smallsetminus\{x_{1}^{W}\}$. It follows that $x_{1}^{W}$ is the only neighbour of $x_{2}^{\prime}$ in $W^{\prime}\smallsetminus U^{\prime}$. This contradicts that $x_{2}$ is a vertex by definition of right-left-shift. ∎
@@
Case: $x_{1}$ is an inward edge. By Lemma 9.33, there is a neighbour $y^{\prime}_{i}$ of $x^{\prime}_{1}$ with $y^{\prime}_{i}=y^{W}_{i}$, and $\pi_{1}$ is not $W$-weird. So $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$ by (6.9.1). Since we assumed earlier that $\pi_{1}$ is weird for some side of $(U,W)$, but showed that $\pi_{1}$ is not $W$-weird, it follows that $\pi_{1}$ is $U$-weird. By Lemma 9.16, there exist two disjoint $\{x^{W}_{1},x^{W}_{2}\}$$\{y^{W}_{1},y^{W}_{2}\}$ paths in $\bigcup_{t\in O-t(0)}G_{t}$. Since $x_{1}$ is an inward edge and $y^{\prime}_{i}=y^{W}_{i}$, the edge $x^{\prime}_{1}y^{\prime}_{i}$ can also be written as $x^{W}_{1}y^{W}_{i}$. This edge together with the two paths forms a set of three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$, so $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $U$ by (6.9.3).
Case: $x_{1}$ is an outward edge. If $\pi_{1}$ is $U$-weird, then we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $e=x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$ by Lemma 9.34. Since $x_{2}\in\pi_{1}$ and $y_{i}^{\prime}\in\hat{\pi}_{2}\cap W$, the edge $e$ is a $\pi_{1}$$(\hat{\pi}_{2}\cap W)$ edge, which witnesses that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $U$ by (6.9.2). So it remains to show that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$. For this, we assume that $\pi_{1}$ is $W$-weird by (6.9.1). By Lemma 9.35, no $y_{i}$ is an inward edge. Since $H_{t(0)}$ is 3-connected, we find three independent $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths in $H_{t(0)}$ by (Menger for independent paths). (6.2). All three paths are subgraphs of $G[U]$ since no $y_{i}$ is an inward edge. Moreover, they are three $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[U]$. This means $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$ by (6.9.3). ∎
@@
By symmetry and by 6.12, it suffices to show the following: if $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird, then there is a $\pi_{1}$$\pi_{2}$ edge in $G$. So we assume that $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird. Then no $y_{i}$ is an inward edge by Lemma 9.33. Without loss of generality, $y_{1}$ is an outward edge. Then by Lemma 9.35, no $x_{i}$ is an inward edge. So we can assume that $x_{1}$ is an outward edge. Then by Lemma 9.34, we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$. Note that $y_{i}$ is a vertex, so $y_{i}^{\prime}=y_{i}\in\pi_{2}$. Then $x_{2}y_{i}^{\prime}$ is the desired edge. ∎
@@
Assume first that $x_{1}$ is an inward edge. Then $y^{\prime}_{1}$ has $\leqslant 1$ neighbour in $W\smallsetminus\{x^{\prime}_{1}\}\supseteq V(G-G_{t(0)})$. But then, by the definition of $(U,W)$ via right-left-shift, $y_{1}$ should be an edge rather than the vertex $y^{\prime}_{1}$, a contradiction.
Assume otherwise that $x_{1}$ is an outward edge. By Lemma 9.38, $x^{W}_{1}\in A$. However, since $y^{\prime}_{1}$ is contained in $A\cap C$ and the Tutte-bagel induces a path-decomposition of $G[A]$, the vertex $y^{\prime}_{1}=y_{1}$ does not neighbour $x^{W}_{1}$, contradicting that $\{x_{1},y_{1}\}$ is $W$-weird. ∎
@@
By Lemma 9.39, neither $\pi_{1}$ nor $\pi_{2}$ is $W$-weird. Hence $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked. It remains to show that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$, say. Since $\pi_{1}$ is not $W$-weird, it remains to consider the case where $\pi_{1}$ is $U$-weird. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$ with the (Three-Paths Lemma). (9.21). Hence $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ via (6.9.3). ∎
@@
If $H_{t}$ is bad, then $(U_{t},W_{t})$ is not even a tetra-separation by Lemma 9.23, let alone totally-nested. So we may assume now that $H_{t}$ is good. In particular, we may assume the (Tutte-bagel torso setting). (9.13) with $t=t(0)$.
By Lemma 9.23, $(U,W)$ is a tetra-separation. It suffices to show that $(U,W)$ is externally 5-connected by 6.16. By Lemma 9.24, $(U,W)$ is half-connected. To see that $(U,W)$ is 3-linked, we combine Lemma 9.30 with Lemma 9.31. By Lemma 9.32, $(U,W)$ is 0-potter-linked. We get that $(U,W)$ is 1-potter-linked and 2-potter-linked from Lemmas 9.36, 9.40 and 9.37. ∎

### v2

@@ 9.5. Totally-nested tetra-separations from good bags
Assume first that $o(B)$ or $o(D)$ has a neighbour in $T$ with a triangle-torso in $\mathcal{O}$. Say this is a neighbour $t$ of $o(B)$. Since $y^{W}_{1}$ is the tip of the triangle $H_{o(B)}$, the vertex $y^{W}_{1}$ is an end of the free edge $f$ of $H_{t}$ by Lemma˜9.15. Let $u$ denote the other end of $f$. Let $v$ denote the tip of $H_{t}$. Note that $\{u,v\}$ is an adhesion-set, so $x^{W}_{1}\in\{u,v\}$. Hence $x^{W}_{1}$ lies in the two triangle-torsos from the potter corner $AD$ and in the triangle-torso $H_{t}$. Thus $o(D)$ and $t$ are neighbours in $T$ by Lemma˜9.15. In particular, $T$ has exactly three nodes. Since $v$ lies in the two triangle-torsos $H_{t}$ and $H_{o(B)}$ while $x^{W}_{1}$ lies in the two triangle-torsos for the potter corner $AD$, we cannot have $v=x^{W}_{1}$ by Lemma˜9.15. Hence $x^{W}_{1}=u$. By Lemma˜9.4, the two adhesion-sets of $H_{t}$ are spanned by edges in $G$. Hence $x^{W}_{1}y^{W}_{1}$ and $x^{W}_{1}vy^{W}_{1}$ are desired paths.
Figure 9.7. Case $i\geqslant j$ in the proof of Lemma˜9.29. Left: $H_{t}$ is a triangle.
@@
Assume otherwise that neither $o(B)$ nor $o(D)$ has a neighbour in $T$ with a triangle-torso. Then the torsos neighbouring $o(B)$ and $o(D)$ via $T$ are 3-connected. Since 3-connected torsos have disjoint adhesion-sets and $i\geqslant j$, we have $|T|=3$. Let $t$ denote the internal node of $T$. Let $\{x^{W}_{1},u\}$ denote the adhesion-set of the edge $o(D)t$, and let $\{y^{W}_{1},v\}$ denote the adhesion-set of the edge $o(B)t$. We find two independent $y^{W}_{1}$$x^{W}_{1}$ paths in $G_{t}$, and are done, unless $H_{t}$ is a 4-wheel with rim $x^{W}_{1}uy^{W}_{1}v$ and $x^{W}_{1}u,y^{W}_{1}v\notin E(G)$, by Lemma˜9.28. But in the latter case, $u$ and $v$ have degree three in $G$, contradicting 4-connectivity. ∎
@@
Say $i=j=1$. If $x_{2}$ or $y_{2}$ is an edge, then we are done by (6.4.1), so assume that both are vertices. By Lemma˜9.25, there is an $x^{\prime}_{1}$$y^{\prime}_{1}$ path $P^{\prime}$ in $G_{t(0)}-\{x^{\prime}_{2},y^{\prime}_{2}\}$. Then $P^{\prime}$ contains a $\hat{x}_{1}$$\hat{y}_{1}$ path $P$ that avoids the vertices $x_{2}$ and $y_{2}$, by Lemma˜9.26. We find two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$ using Lemma˜9.29. Hence $\{x_{1},y_{1}\}$ is 3-linked around $S(U,W)$ by (6.4.3). ∎
@@
If $y_{1}$ or $y_{2}$ is an edge, we are done, so assume that both are vertices. There exist two independent $x^{W}_{1}$$x^{W}_{2}$ paths in $G[W]-\{y_{1},y_{2}\}$ by the (Three-Paths Lemma). (9.21). So it remains to find an $x^{\prime}_{1}$$x^{\prime}_{2}$ path in $G^{\prime}:=G_{t_{0}}-\{y^{\prime}_{1},y^{\prime}_{2}\}$ (as any such path contains a suitable $\hat{x}_{1}$$\hat{x}_{2}$ path by Lemma˜9.26). Assume for a contradiction that such a path is missing. Then $H_{t(0)}$ is a $K_{4}$ by Lemma˜9.17. Since both $y_{1}$ and $y_{2}$ are vertices, the torso $H_{t(1)}$ is 3-connected by Lemma˜9.14. Since $H_{t(0)}$ is good, it follows that $H_{t(-1)}$ is a 4-cycle. So $x_{1}$ and $x_{2}$ are the two free edges on this 4-cycle. By 4-connectivity of $G$, both $x^{\prime}_{1}$ and $x^{\prime}_{2}$ have degree at least four. Since $H_{t(-1)}$ is a 4-cycle and $H_{t(0)}$ is a $K_{4}$, it follows that the edge $x^{\prime}_{1}x^{\prime}_{2}$ must exist in $G$, contradicting that there is no $x^{\prime}_{1}$$x^{\prime}_{2}$ path in $G^{\prime}:=G_{t_{0}}-\{y^{\prime}_{1},y^{\prime}_{2}\}$. ∎
@@
Suppose first that $\pi_{1}=\{x_{1},x_{2}\}$ and $\pi_{2}=\{y_{1},y_{2}\}$. We find two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{1},P_{2}$ in $G[W]$ using Lemmas˜9.16 and 9.14. The torso $H_{t(0)}$ is 3-connected by assumption, and hence it contains three independent $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths by (Menger for independent paths). (6.2). None of these paths use torso-edges, so all three paths are subgraphs of the bag $G_{t(0)}$. The three $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths can be shortened to three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $Q_{1},Q_{2},Q_{3}$ by Lemma˜9.26. Hence $\pi_{1}$ is 0-potter-linked to $\pi_{2}$.
Suppose otherwise that $\pi_{1}=\{x_{1},y_{1}\}$ and $\pi_{2}=\{x_{2},y_{2}\}$, say. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{1},P_{2},P_{3}$ in $G[W]-x^{\prime}_{1}x^{\prime}_{2}-y^{\prime}_{1}y^{\prime}_{2}$ using the (Three-Paths Lemma). (9.21). By Lemma˜9.28, we find two independent $\{x^{\prime}_{1},y^{\prime}_{1}\}$$\{x^{\prime}_{2},y^{\prime}_{2}\}$ paths in the bag $G_{t(0)}$, which shorten to two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{4},P_{5}$ in $G_{t(0)}$ by Lemma˜9.26, unless $H_{t(0)}$ is a 4-wheel and both torso-edges $x^{\prime}_{1}x^{\prime}_{2},y^{\prime}_{1}y^{\prime}_{2}$ are missing in $G$. But the latter is impossible: since $H_{t(0)}$ is good, some neighbouring torso would be a triangle or a 4-cycle, but then a vertex in the adhesion-set would have degree three in $G$, contradicting 4-connectivity. ∎
@@
˜(i). Since $H_{t(0)}$ is 3-connected, the vertex $x^{\prime}_{1}$ has $\geqslant 3$ neighbours in $H_{t(0)}$. Let $u$ be a neighbour of $x^{\prime}_{1}$ in $H_{t(0)}$ distinct from $x^{\prime}_{2}$ and the end of the edge $x_{1}$ besides $x^{\prime}_{1}$. All interior vertices of $H_{t(0)}$ are contained in $U\smallsetminus W$, so $u\in\{y^{\prime}_{1},y^{\prime}_{2}\}$. Say $u=y^{\prime}_{1}$. Assume for a contradiction that $y^{W}_{1}\notin H_{t(0)}$. Then $x^{\prime}_{1}y^{\prime}_{1}$ is a $(W\smallsetminus U)$$(U\smallsetminus W)$ edge, so $(U,W)$ violates the matching-condition, a contradiction.
˜(ii). Assume for a contradiction that $\{x_{1},x_{2}\}$ is $W$-weird. Then $x_{2}=x^{\prime}_{2}$ since $x_{1}$ is an edge by assumption. Moreover, $x^{\prime}_{2}$ has exactly one neighbour outside $H_{t(0)}$ by weirdness. But then $x_{2}$ should be an edge in $S(U,W)$ by the definition of right-left-reduction of $(U^{\prime},W^{\prime})$, contradicting that $x_{2}$ is a vertex. ∎
@@
Assume otherwise, say $y_{1}$ is an inward edge. Then by Lemma˜9.33 we have $y^{W}_{1}=y^{\prime}_{1}$ and some $x^{\prime}_{j}$ is a neighbour of $y^{W}_{1}$. Then $j\neq 1$ since otherwise $x^{\prime}_{1}y^{W}_{1}$ is a $(U\smallsetminus W)$$(W\smallsetminus U)$ edge, contradicting the matching-condition. So $j=2$. Then $y_{1}^{\prime}=y_{1}^{W}$ is the only neighbour of $x_{2}$ in $W\smallsetminus\{x_{1}^{W}\}$. It follows that $x_{1}^{W}$ is the only neighbour of $x_{2}^{\prime}$ in $W^{\prime}\smallsetminus U^{\prime}$. This contradicts that $x_{2}$ is a vertex by definition of right-left-reduction. ∎
@@
Case: $x_{1}$ is an inward edge. By Lemma˜9.33, there is a neighbour $y^{\prime}_{i}$ of $x^{\prime}_{1}$ with $y^{\prime}_{i}=y^{W}_{i}$, and $\pi_{1}$ is not $W$-weird. So $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$ by (6.9.1). Since we assumed earlier that $\pi_{1}$ is weird for some side of $(U,W)$, but showed that $\pi_{1}$ is not $W$-weird, it follows that $\pi_{1}$ is $U$-weird. By Lemma˜9.16, there exist two disjoint $\{x^{W}_{1},x^{W}_{2}\}$$\{y^{W}_{1},y^{W}_{2}\}$ paths in $\bigcup_{t\in O-t(0)}G_{t}$. Since $x_{1}$ is an inward edge and $y^{\prime}_{i}=y^{W}_{i}$, the edge $x^{\prime}_{1}y^{\prime}_{i}$ can also be written as $x^{W}_{1}y^{W}_{i}$. This edge together with the two paths forms a set of three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$, so $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $U$ by (6.9.3).
Case: $x_{1}$ is an outward edge. If $\pi_{1}$ is $U$-weird, then we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $e=x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$ by Lemma˜9.34. Since $x_{2}\in\pi_{1}$ and $y_{i}^{\prime}\in\hat{\pi}_{2}\cap W$, the edge $e$ is a $\pi_{1}$$(\hat{\pi}_{2}\cap W)$ edge, which witnesses that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $U$ by (6.9.2). So it remains to show that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$. For this, we assume that $\pi_{1}$ is $W$-weird by (6.9.1). By Lemma˜9.35, no $y_{i}$ is an inward edge. Since $H_{t(0)}$ is 3-connected, we find three independent $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths in $H_{t(0)}$ by (Menger for independent paths). (6.2). All three paths are subgraphs of $G[U]$ since no $y_{i}$ is an inward edge. Moreover, they are three $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[U]$. This means $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$ by (6.9.3). ∎
@@
By symmetry and by ˜6.12, it suffices to show the following: if $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird, then there is a $\pi_{1}$$\pi_{2}$ edge in $G$. So we assume that $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird. Then no $y_{i}$ is an inward edge by Lemma˜9.33. Without loss of generality, $y_{1}$ is an outward edge. Then by Lemma˜9.35, no $x_{i}$ is an inward edge. So we can assume that $x_{1}$ is an outward edge. Then by Lemma˜9.34, we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$. Note that $y_{i}$ is a vertex, so $y_{i}^{\prime}=y_{i}\in\pi_{2}$. Then $x_{2}y_{i}^{\prime}$ is the desired edge. ∎
@@
Assume first that $x_{1}$ is an inward edge. Then $y^{\prime}_{1}$ has $\leqslant 1$ neighbour in $W\smallsetminus\{x^{\prime}_{1}\}\supseteq V(G-G_{t(0)})$. But then, by the definition of $(U,W)$ via right-left-reduction, $y_{1}$ should be an edge rather than the vertex $y^{\prime}_{1}$, a contradiction.
Assume otherwise that $x_{1}$ is an outward edge. By Lemma˜9.38, $x^{W}_{1}\in A$. However, since $y^{\prime}_{1}$ is contained in $A\cap C$ and the Tutte-bagel induces a path-decomposition of $G[A]$, the vertex $y^{\prime}_{1}=y_{1}$ does not neighbour $x^{W}_{1}$, contradicting that $\{x_{1},y_{1}\}$ is $W$-weird. ∎
@@
By Lemma˜9.39, neither $\pi_{1}$ nor $\pi_{2}$ is $W$-weird. Hence $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked. It remains to show that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$, say. Since $\pi_{1}$ is not $W$-weird, it remains to consider the case where $\pi_{1}$ is $U$-weird. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$ with the (Three-Paths Lemma). (9.21). Hence $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ via (6.9.3). ∎
@@
If $H_{t}$ is bad, then $(U_{t},W_{t})$ is not even a tetra-separation by Lemma˜9.23, let alone totally-nested. So we may assume now that $H_{t}$ is good. In particular, we may assume the (Tutte-bagel torso setting). (9.13) with $t=t(0)$.
By Lemma˜9.23, $(U,W)$ is a tetra-separation. It suffices to show that $(U,W)$ is externally 5-connected by ˜6.16. By Lemma˜9.24, $(U,W)$ is half-connected. To see that $(U,W)$ is 3-linked, we combine Lemma˜9.30 with Lemma˜9.31. By Lemma˜9.32, $(U,W)$ is 0-potter-linked. We get that $(U,W)$ is 1-potter-linked and 2-potter-linked from Lemmas˜9.36, 9.40 and 9.37. ∎